The auto repair shop of Quality Motor Company uses standards to control the labor time and labor cost in the shop. The standard labor cost for a motor tune-up is given below:

  Job Standard Hours   Standard Rate   Standard Cost  
  Motor tune-up 2.10         $6.50         $13.65       

The record showing the time spent in the shop last week on motor tune-ups has been misplaced. However, the shop supervisor recalls that 140 tune-ups were completed during the week, and the controller recalls the following variance data relating to tune-ups:

  Labor rate variance $ 867 F
  Labor spending variance $ 537 U
Required:
1. Determine the number of actual labor-hours spent on tune-ups during the week.
2.

Determine the actual hourly rate of pay for tune-ups last week.
Answer #1

Labor efficiency variance = Labor spending variance - Labor rate variance

= 537U - 867F

= 1404 U

Labor efficiency variance = (SH-AH) *SR

(2.10*140 - AH) *6.50 = - 1404

(294 - AH) = - 216

AH = 294 + 216 = 510

Labor Rate Variance = (SR-AR) *AH

(6.50 - AR) *510 = 867

(6.50 - AR) = 1.70

AR = 4.80

Actual Labor Hours 510 hours
Actual hourly rate 4.80 per hour